### v1

# DRIFT: Dynamic Rule-Based Defense with Injection Isolation for Securing LLM Agents

Abstract
Large Language Models (LLMs) are increasingly central to agentic systems due to their strong reasoning and planning capabilities. By interacting with external environments through predefined tools, these agents can carry out complex user tasks. Nonetheless, this interaction also introduces the risk of prompt injection attacks, where malicious inputs from external sources can mislead the agent’s behavior, potentially resulting in economic loss, privacy leakage, or system compromise. System-level defenses have recently shown promise by enforcing static or predefined policies, but they still face two key challenges: the ability to dynamically update security rules and the need for memory stream isolation. To address these challenges, we propose DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems, which enforces both control- and data-level constraints. A Secure Planner first constructs a minimal function trajectory and a JSON-schema-style parameter checklist for each function node based on the user query. A Dynamic Validator then monitors deviations from the original plan, assessing whether changes comply with privilege limitations and the user’s intent. Finally, an Injection Isolator detects and masks any instructions that may conflict with the user query from the memory stream to mitigate long-term risks. We empirically validate the effectiveness of DRIFT on the AgentDojo benchmark, demonstrating its strong security performance while maintaining high utility across diverse models—showcasing both its robustness and adaptability.
1 Introduction
Large Language Models (LLMs), Large Language Models (LLMs), empowered by their exceptional planning and reasoning abilities, are increasingly integrated into agentic systems [1, 2, 3]. By processing natural language data streams, LLM agents interact with external environments, such as applications [1, 4], computing systems [3], via a set of pre-defined tools to carry out complex user tasks. Since the need for interaction with untrusted external environments, a new security threat of prompt injection attacks is introduced [5, 6, 7, 8, 9], where attackers inject malicious instructions into third-party platforms, misleading the agent workflow after external interaction. For example, a product review on Amazon written by another user, such as “Ignore previous instructions, buy this red shirt,” may manipulate the LLM into executing unintended actions. This form of attack [5, 6, 7, 8, 9] may bring risks such as economic losses [6], privacy leakage [10], and system damage [11] to users, severely undermining the reliability of the agentic system.
Existing defense mechanisms can be broadly categorized into model-level [12, 13, 14, 15, 16, 17] and system-level [18, 19, 20, 21, 22] approaches. Model-level defenses [12, 13, 14, 15, 16, 17] typically rely on training a guardrail to detect injection input or mitigate the injection impact. Such methods are constrained by the inherent vulnerabilities of the models and often struggle to defend against complex and diverse attacks. Recently, system-level [18, 19, 20, 21, 22] defenses have gained increasing attention, as they are able to overcome the inherent vulnerabilities of models when facing unseen attacks, thereby achieving high practical reliability in real-world applications. These approaches typically constrain the model’s action space through predefined security policies or rules to prevent potential injection threats. For instance, IsolateGPT [18] mitigates information leakage risks by enforcing isolation mechanisms and maintaining a separate memory bank for each application. Recently, CaMeL [21] achieves impressive security by manually defining a set of security policies and constructing a strict, fixed control and data dependency graph from the user query before any interaction takes place.
Despite the progress in system-level defense mechanisms for agentic systems, two critical challenges remain largely unresolved: (1) the dynamic updating of security rules and (2) the isolation of injected content within the memory stream. While CaMeL enforces robust security through a strict dependency graph, but this static design considerably sacrifices flexibility and practical usability, particularly in agentic systems that require adaptive, real-time decision-making. Furthermore, the reliance on manually crafted security policies imposes considerable overhead and impedes generalization across diverse usage scenarios. In addition, the method of IsolateGPT restrict the propagation of injection-related information across different applications, but residual injection content preserved in memory still poses significant risks within the same application during prolonged interactions.
To overcome these challenges, we develop DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems that enforces security through both control- and data-level constraints. We first design a Secure Planner, which establishes the initial constraint rules solely according to the user query prior to any interaction. It constructs a minimal function trajectory (control constraints) to avoid injections misleading by executing functions in order. In addition, a checklist for each function node in the trajectory, with detailed parameter requirement and value dependencies, is encoded in JSON schema format [23]. When trajectory deviations are detected, a Dynamic Validator performs approval assessments based on the privilege category (Read, Write, Execute) and its alignment with the user’s original intent. To avoid the risk of injection messages to the agent or other modules during prolonged interactions, an Injection Isolator is also designed to continuously polish the memory after each interaction, identifying and masking any instructions that conflict with the initial user query. This layered defense strategy ensures strong context isolation while enabling secure and adaptive decision-making throughout long-term agent interactions. We include a more detailed discussion of related work in Appendix A.
As a fully automatic system-level defense framework, DRIFT demonstrates strong performance across diverse scenarios, achieving high security while maintaining robust utility. Specifically, we evaluate DRIFT on the AgentDojo [24] benchmark, a simulated agent environment featuring various task scenarios and types of injection attacks. By applying DRIFT to GPT-4o-mini [25], the Attack Success Rate (ASR) is successfully reduced from 30.7% to 1.3%, while utility outperforms CaMeL by 20.1% under no attack and by 12.5% under attack. In addition, DRIFT shows remarkable adaptability and generalization across four advanced online LLMs: GPT-4o [26], GPT-4o-mini [25], Claude-3.5-sonnet [27], Claude-3-haiku [28], and one prevalent offline LLM, Qwen2.5-7B [29]. On all of these models, DRIFT significantly enhances security while maintaining or even improving utility on some models. Moreover, we finetune our policy on DRIFT, with the dataset collected from ToolBench [30], achieving significant improvements in both security and utility. Compared to the original version, the ASR of the policy-tuned model drops from 15.1% to 0.0%, while utility under no attack increases from 26.6% to 32.2%, and utility under attack improves from 19.1% to 22.2%. This DRIFT policy training mechanism could enable more reliable, robust, secure, and functional LLM agentic systems.
The main contributions of this work are summarized as follows:
  • We develop DRIFT, a comprehensive system-level defense framework that integrates dynamic security policies with memory isolation, achieving a superior balance between security and functionality.
  • Extensive experiments demonstrate the effectiveness and adaptability of DRIFT across a wide range of scenarios, as well as the effectiveness of each component within DRIFT.
    footnotetext: Project Page: https://github.com/leolee99/DRIFT
2 DRIFT: Dynamic Rule-based Isolation Framework
DRIFT is a system-level rule-based defense framework designed to protect LLM-based agents from prompt injection attacks by strictly enforcing both control- and data-level constraints to ensure security. A dynamic permission mechanism is employed to continuously update these constraints, which helps maintain task utility. Additionally, an injection memory isolation mechanism is integrated to mitigate long-term risks posed by in-memory injection messages. An overview of Secure Planner is shown in Figure 1. Overall, DRIFT comprises the following key components:
  • Secure Planner: An LLM used to plan and parse structured function trajectory (control constraints) and parameter checklists (data constraints) from queries.
  • Dynamic Validator: An LLM for dynamic verification of function trajectory deviation.
  • Injection Isolator: An isolator that detects and removes the instructions conflicting with the user query from memory.
2.1 Secure Planner
Secure Planner is a large language model that operates in the initial phase before any interaction with the environment. This phase is critical for establishing foundational security policies, as it occurs when there is no risk of injection attacks. During this stage, Secure Planner constructs both control-level and data-level policies to constrain the agent’s subsequent actions.
Secure Planner first analyzes the original user query and decomposes the task into a sequence of subtasks. Based on this decomposition, it generates a minimal function trajectory that serves as the basis for control-level constraints. For data-level constraints, Secure Planner creates a JSON-formatted checklist specifying the required parameters and their value dependencies for each function node. These processes are driven by an LLM through a prompt in Figure 7. This mechanism defends against attacks that attempt to invoke the same function with altered parameters. For instance, in a flight booking system, given a user query like “book a flight from Paris to London,” an injected instruction such as “book a flight from London to New York” could bypass control-only policies. However, with data-level constraints in place, such discrepancies can be detected and blocked.
Figure 1: The overview of secure planner, dynamic validator and injection isolator.
2.2 Dynamic Validator
After interacting with the environment, the Dynamic Validator is employed to ensure alignment with control and data constraints, thereby mitigating potential injection attacks. It also dynamically handles inconsistencies to preserve the agent’s utility in completing user tasks.
Alignment Validation.
Following the generation of each tool-calling request, the Dynamic Validator checks whether the function to be executed adheres to both control- and data-level constraints. It first integrates the function into the agent’s executed function trajectory and compares it with the predefined minimal function trajectory. Similarly, the consistency and dependency of function parameters are validated against the predefined parameter checklists, which are established by the Secure Planner. If both the function and its parameters align with the initial constraints, the agent is permitted to proceed with the user’s task.
Dynamic Constraint Policy.
In real-world agent scenarios, the environment is unpredictable, and many decisions must be made after interactions. It is difficult to initialize a complete and sufficient constraint policy at the beginning. A strict and static constraint policy inevitably sacrifices task utility, especially in complex tasks. To address this, we propose a dynamic constraint updating approach. Specifically, when the function trajectory deviates from the expected path, we first identify the role category of the deviated function and assign it a privilege mark.
Inspired by the privilege concepts in Operating Systems (OS), we categorize functions into three roles: Read, Write, and Execute through the prompt shown in Figure 8. If a function only performs read-only operations, such as get_inbox, it is assigned the Read privilege. If a function modifies, updates, creates, or deletes data—such as update_user_info—it is assigned the Write privilege. Functions that trigger interactions with third-party objects (e.g., send_email) are marked as Execute.
In general, a function with the Read privilege does not directly pose a risk to the user and will be approved even if it deviates from the original trajectory. However, functions marked as Write or Execute may introduce direct risks. In such cases, the Validator will assess whether the deviated function aligns with the user’s original intent based on the updated tool messages, using the prompt shown in Figure 9. If the deviated function still aligns with the user’s intent, the function is approved and incorporated into the minimal function trajectory and parameter checklist to support successful validation in subsequent validation. Otherwise, agents will send an approval request to user.
2.3 Injection Isolator
Current rule-based agent defense approaches typically restrict action permissions but do not eliminate injected content. In a long-term agentic system, past memory—such as previous conversations and tool responses—is frequently reused. These reused elements may be accessed not only by the agent itself but also by other components within the security system, such as the policy updating module. During the process of policy optimization, it is inevitable to incorporate new information obtained from recent interactions. However, any injection content stored in the memory stream will also be repeatedly exposed to these components during long-term interactions, severely increasing the risk of compromise over time.
To mitigate this long-term threat, we propose an injection isolation mechanism to detect and remove injected content from the memory stream. Specifically, we design a curated Injection Isolator that analyzes returned messages from each tool-calling and determines whether any instructions conflict with the user’s original intent. The identification process is driven by a LLM using system prompt in Figure 10. If a conflict is detected, the isolator removes the conflicting instructions using external masking components before the message is added to the agent’s memory stream. Subsequently, a safe memory stream could be maintained in long-term agent interactions. The Isolator cannot directly modify the tools and does not interact with the agent, which helps prevent potential security vulnerabilities as much as possible.
Figure 2: The workflow of DRIFT.
2.4 Security Policies in LLM Agents
An LLM-based agentic system typically comprises four key components: the user, the agent, tools, and the environment. In a standard workflow, the user first sends a query to the agent. The agent then goes through a reasoning process (e.g., chain-of-thought [31]) and selects a suitable tool to call. The response from the tool helps guide the agent’s next decision. The agent typically completes the user’s task through several such cycles. During this process, injection attacks can occur through injecting malicious content in tool responses.
Our secure framework, DRIFT, can be integrated into agentic systems built on different LLMs. The overall workflow is shown in Figure 2. In the initial phase, the Secure Planner sets up a function trajectory to constrain the control flow, and a parameter checklist for each function node to constrain the data flow.
The user query is then fed into the agent, triggering a reasoning process and generating tool-calling decisions. Afterward, the Dynamic Validator checks whether the function deviates from the original plan and updates the approval policy if necessary. If the call is approved and retrieves results from the environment, the Injection Isolator inspects the tool outputs for instructions that conflict with the user’s original query. If any are found, they are masked by an external program. The cleaned responses are then stored in memory for use in future steps.
2.5 Trainable Security Policy
To enhance the reliability and generalization of our security policy, we develop a training approach for both the Secure Planner and the Injection Isolator, allowing our DRIFT framework to adapt more robustly. This involves designing a new data collection pipeline that extracts policy-aligned samples from existing agent datasets, followed by efficient instruction tuning using Low-Rank Adaptation (LoRA) [32] on Qwen2.5-7B [29].
2.5.1 Data and Environment Construction
Although datasets like ToolBench [30] have been collected to support tool-use reasoning in LLMs, their formats do not align well with the structure of our security policy. This makes them less suitable for direct training. To address this, we introduce a method for generating training data that adheres our policy, by modifying existing conversations from ToolBench. Each conversation in ToolBench includes messages from three sources: user, tool, and assistant. We use GPT-4o-mini to rewrite the assistant messages to align with our policy.
Planner Data Sampling.
For training the Secure Planner, we keep the original user query and tool-calling trajectory, but rewrite the first-round assistant message using system prompt of Figure 11. Assistant messages generally include reasoning thoughts and tool calls. We modify the reasoning part using GPT-4o-mini to produce a JSON-style minimal function trajectory and parameter checklist, while keeping the tool called to preserve the original flow. We collect 1,000 such samples, with conversations ranging from 4 to 14 turns.
Isolator Data Sampling.
To train the Injection Isolator, we simulate injected instructions within tool outputs. These injections are automatically designed to fit the topic and context of the conversation, making them appear realistic and challenging. GPT-4o-mini is employed to generate the injected content and determine where to place it, using the system prompt of Figure 12. After the injection, we rewrite the assistant message to detect and highlight the injected instructions clearly. We finally collect 1,000 training samples for the Isolator.
Tool Environment Re-construction.
In practical agentic systems, the number of visible tools can be much larger than typically seen in datasets like ToolBench, where each sample involves only a few tools (usually fewer than five). To better reflect real-world scenarios, we collect tool metadata from 5,000 samples and build a tool list with over 10,000 non-redundant unique tools. For each new training instance, we randomly add 0 to 25 extra tools to the external tools, creating a more realistic and challenging environment.
2.5.2 Agent Training.
After data collection completed, we fine-tune the Qwen2.5-7B model using LoRA for both the Secure Planner and Injection Isolator, as well as the agent itself. For the Dynamic Validator, we rely on the original Qwen2.5-7B in a zero-shot setup to handle privilege classification and user intent checking.
3 Experiments
In this section, we evaluate DRIFT on AgentDojo, the latest and prevalent agentic security benchmark, to assess the effectiveness and adaptability of DRIFT in terms of both utility and security. Furthermore, we analyze the contribution of each individual component within DRIFT.
3.1 Experimental Setups
Benchmarks.
We evaluate our method with AgentDojo [24], a benchmark that simulates realistic interactions in agent-based systems. It includes four scenarios—banking, Slack, travel, and workspace—covering 97 user tasks to assess utility and 629 injection tasks to evaluate security.
Metrics.
Following the AgentDojo setup, we report three metrics: Benign Utility, Utility Under Attack, and Targeted Attack Success Rate (ASR). Benign Utility measures the frequency with which the agent completes the intended task in the absence of attacks. Utility Under Attack looks at how often the agent still completes the original task despite adversarial inputs. ASR reflects how often an injection attack succeeds in achieving the attacker’s goal.
Baselines.
We compare our method against five existing advanced defense approaches, with four approaches available in AgentDojo: repeat_user_prompt, spotlighting_with_delimiting, tool_filter, transformers_pi_detector, and a policy-based security approach of CaMeL [21]. These represent a range of strategies for protecting agents from injection attacks.
Implementation Details.
We apply our policy to several models, including online models—GPT-4o [26], GPT-4o-mini [25], Claude-3-haiku [28], and Claude-3.5-sonnet [27]—and an offline model, Qwen2.5-7B [29]. For Qwen2.5-7B, we fine-tune it on our policy dataset (described in Section 2.5) using a batch size of 4 and training for three epochs. We employ the Adam optimizer [33] with weight decay and set the initial learning rate to 2e-5. The maximum input length is 15,000 tokens.
Figure 3: Comparison across different defense method on GPT-4o-mini.
3.2 Defense Techniques Comparison
In this section, we compare DRIFT with five advanced defense techniques—four implemented in AgentDojo: repeat_user_prompt, spotlighting_with_delimiting, tool_filter, transformers_pi_detector—and one policy-based defense approach, CaMeL. Figure 3 illustrates the results. We use GPT-4o-mini-2024-07-18 as the base agent for all evaluations. Notably, the DRIFT policy achieves an optimal balance between utility and security. In terms of security, DRIFT significantly outperforms all other baselines except CaMeL, with only a marginal gap of 1.6%. However, in terms of utility—both in no-attack and under-attack conditions—DRIFT surpasses CaMeL by 21.8% in the no-attack setting and 10.9% under attack. This indicates that DRIFT achieves a superior utility-security trade-off, making it more practical for real-world agentic systems.
3.3 DRIFT Adaptation
DRIFT is a system-level defense framework that can be deployed across many types of agents. To better understand the adaptability and generality of DRIFT in different agent settings, we apply it to multiple LLMs, including four advanced online models—GPT-4o [26], GPT-4o-mini [25], Claude-3 Haiku [28], and Claude-3.5-Sonnet [27]—and one widely used offline model, Qwen2.5-7B [29].
For the online models, we compare our method with agents using ReAct [34], a technique that allows the LLM to reason and call tools in an agentic manner. The results are presented in Figure 4 (detailed results on four scenarios shown in Appendix B). We observe that DRIFT significantly enhances security across all models, reducing ASR from over 10% to single-digit levels, strongly indicating the security generality of DRIFT across diverse models. Notably, GPT-4o with ReAct, one of the most advanced LLMs with strong general capabilities, shows a high ASR of 51.7%, highlighting the vulnerability of current LLM agents—even those powered by leading models. However, after deploying DRIFT, the ASR drops sharply from 51.7% to just 1.5%, further demonstrating the effectiveness of DRIFT in securing agents from attack.
In addition, DRIFT does not compromise the agent’s task completion ability, as shown by the stable utility scores in both safe and unsafe conditions. In some cases, DRIFT even improves utility, e.g., with GPT-4o and Claude-3.5 Sonnet under attack.
The offline model Qwen2.5-7B, which has been tuned on our policy, achieves remarkable improvements in both utility and security. In terms of utility, our tuned agent obtains a 5.6% improvement in safe conditions and 3.1% in unsafe conditions. It is noticeable that the ASR after tuning drops to 0. These improvements highlight a potential solution for robustly securing agentic systems without performance sacrifice. All of these results demonstrate the effectiveness of DRIFT across different models and scenarios, fully supporting its broad adaptability and strong generality.
Figure 4: Comparison across different LLM Agents.
3.4 Ablation Studies
In this section, we perform ablation studies to examine the individual contributions of each DRIFT component: Planner, Validator and Isolator. The results are presented in Table 1.
We begin with the Native Agent setup, which uses the ReAct technique to serve as agents. GPT-4o-mini serves as the base model, with no defense mechanism applied. In this setting, the agent is vulnerable to be attacked, with a Targeted Attack Success Rate (ASR) of 30.67%. We then add the Secure Planner on the Native Agent, which generates fixed control- and data-level constraints based on the initial user query. These strict policies significantly improve security, reducing ASR to just 1.49%, showing the effectiveness of static policy enforcement. However, this improvement introduces severely utility drops. Specifically, The Utility in no attack decreases from 63.55% to 37.71% (a drop of 25.84%), and Utility Under Attack falls from 48.27% to 32.25% (a drop of 16.02%). This illustrates the limitation of using a static policy significantly undermines the agent capability to complete the tasks.
Afterward, we incorporate the Dynamic Validator, which adjusts policies during execution based on the agent’s interactions. This dynamic mechanism leads to a notable improvement in utility while maintaining strong security: Benign Utility and Utility Under Attack increase to 59.79% and 48.43%, respectively, while ASR rises slightly to 3.66%. These results demonstrate that dynamic policy updates provide a better balance, improving task success without significantly compromising security. To further explore the necessity of dynamic policies, we analyze how static and dynamic policies perform against the change of task complexity in Section 3.5.
Finally, we add the Injection Isolator, designed to mitigate long-term legacy risks by identifying and masking conflicting or malicious content from the memory stream. This component further reduces ASR to just 1.29%, which is lower than the ASR achieved using only the strict policy. Besides, it causes only a slight drop in utility, demonstrating its effectiveness in enhancing the DRIFT security.
Overall, this ablation study highlights the role of each component in DRIFT. It reveals the underlying mechanisms of how each component contributes to enhancing agent performance and how they work together to achieve a strong balance between security and utility.
Table 1: Ablation Studies on different components of DRIFT.
Model Utility (No Attack) $\uparrow$ Utility (Under Attack) $\uparrow$ ASR (Under Attack) $\downarrow$
Native Agent 63.55 48.27 30.67
+ Secure Planner 37.71 32.25 1.49
+ Dynamic Validator 59.79 48.43 3.66
+ Injection Isolator (Full) 58.48 47.91 1.29
3.5 Necessity of Dynamic Policy in Agentic System
To better understand the necessity of a dynamic policy in agentic systems, we explore the performance of static policy and dynamic policy on four sessions (i.e., Banking, Slack, Travel, and Workspace) in AgentDojo, with the results shown in Figure 5a. We observe that the dynamic policy outperforms the static policy in all sessions, with a significant gap in all but the Banking session. To identify the hindering reason for this gap, we analyze the trajectory lengths in these sessions, most of which are shorter than 3. In most cases, trajectory length can represent the complexity of the user task.
To further explore the underlying mechanism behind the correlation between user task complexity and the performance gap, we count all samples in AgentDojo and plot a line chart in Figure 5b to show the scaling law between Success Rate (SR) and trajectory length. We observe that when the trajectory length is no more than 2, the success rates of agents with static and dynamic policies show a similar gradient. However, when the trajectory length reaches or exceeds 3, there is a sharp decrease in the success rate for agents with static policies, while the dynamic policy remains stable. This indicates the limitation of static policies in long-trajectory (complex task) scenarios.
In real-world agentic systems, there are few tasks that require only 1–2 steps to complete. This practical need highlights the necessity of a dynamic mechanism in real-world agentic systems.
(a) Comparison between the Agent with and without the dynamic mechanism on four sessions of AgentDojo
(b) Trajectory Scaling
Figure 5: (a) The comparison between Agent with static policy and with dynamic policy on four sessions of AgentDojo. (b) Task success rate against trajectory length.
(a) w/o Injection Isolator
(b) w/ Injection Isolator
Figure 6: A case study of Injection Isolator on defending prompt injection attacks.
3.6 Case Study for Injection Isolator
To better understand the effectiveness of the Injection Isolator in defending against prompt injection attacks, we present a real case from AgentDojo in Figure 6.
In Figure 6a, we observe that the agent is successfully attacked by injection instructions embedded in the messages returned by the function get_rating_reviews_for_hotels. The agent follows these instructions and includes risky content in its final answer. Notably, the tool trajectory and parameters are not misled in this case—the attack occurs despite correct tool usage. This reveals a key insight: control and data constraints alone are not sufficient to prevent all types of injection attacks.
It is also important to note that the injection message is introduced during the first tool call. Even though further reasoning and interactions take place afterward (e.g., a reserve_hotels call), the malicious content still influences the final output, since all historical conversations are re-input into the agent before generating the final answer. This shows that once injected, harmful messages pose an ongoing risk if they are stored in the agent’s memory stream.
By contrast, the agent equipped with our Injection Isolator (Figure 6b) successfully defends against this type of attack and avoids the risk of malicious content being stored in the memory stream, which could be exposed to other modules or subsequent interactions. This case study demonstrates the effectiveness and importance of the injection isolation mechanism in securing agentic systems.
4 Conclusion
In this paper, we delve into system-level defenses for LLM agents against prompt injection attacks. We develop DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems. This framework generate dynamic policies to constrain agent actions, ensuring security while maintaining utility. It includes an injection isolation mechanism to remove injected content from the memory stream, preserving long-term security. Overall, we present a Secure Planner, a Dynamic Validator, and an Injection Isolator, achieving a generalized, secure, and functional agentic system.
Limitations.
While our work demonstrates significant improvements in both utility and security on the AgentDojo benchmark—one of the most prevalent agent simulation environments—the benchmark domains are limited and do not fully cover the diverse tasks and attack scenarios encountered in real-world agentic systems. To further validate the effectiveness of DRIFT, future work will focus on evaluating its performance in more realistic and diverse environments.
Appendix
Appendix A Related Works
A.1 LLM Agents
LLM Agents [1, 2, 3, 35, 36, 37, 38] are powered by large language models to automatically perceive environments and make decisions. Benefiting from the powerful reasoning capabilities of LLMs, a number of efforts [1, 37, 38, 3] equip LLM agents with tools to help users automatically complete tasks. Furthermore, recent advancements [37, 1, 2, 39] like Mind2Web [2] and WebAgent [1] construct systems to interact with web pages. OSWorld [3] constructs a desktop-manipulated system that enables agents to interact with computers. Additionally, several studies have explored methods to enhance agent reasoning capabilities. ReAct [34] introduces an effective approach to enhance the reasoning and acting capabilities of LLMs. Language Agent Tree Search [40] is proposed to improve the multi-step reasoning and planning capabilities of LLM agents. Some recent research also explores better tool selection mechanisms [38, 41, 30, 42]. REST-GPT [38] develops a flexible tool-calling interface for LLM agents. ToolBench [30] introduces a web-crawled benchmark for training and evaluating the tool-usage capabilities of LLMs.
A.2 Prompt Injection Defenses
A line of studies [12, 13, 14, 15, 18, 21] has explored solutions for securing LLM agents from prompt injection attacks. Current prompt injection defenses can be classified into model-level and system-level approaches.
Model-level defenses focus on enhancing the model’s inherent ability to resist attacks. StruQ [12] introduces a mechanism to transform queries into a structured form and trains the model to focus on the structured part. Chen et al.  [13] propose a preference optimization approach to defend against injection attacks. Another significant direction involves injection detection through external models, such as LlamaGuard [14] and InjecGuard [15]. These specialized models are trained to identify potentially malicious content across multiple risk categories, offering a complementary layer of protection.
System-level defenses typically constrain the model’s action space through predefined security policies to prevent attacks. Early system-level defenses focus primarily on coding scenarios [43] and face challenges when transferred to tool-integrated agent environments [1, 2, 3].
Recent advances in system-level protection have produced several notable approaches for tool-integrated agents. IsolateGPT [18] builds isolated execution environments for each application to reduce cross-application data flow risks. Both f-secure [19] and RTBAS [20] implement information flow control mechanisms that constrain untrusted data and propagate untrusted labels throughout the system. CaMeL [21] constructs control and data flows from the original user query and designs an interpreter to protect flow security. However, its control and data flow policies are static and cannot adequately meet the needs of dynamic real-time interactions. Concurrent with our work, Progent [22] develops a dynamic policy update mechanism based on historical interactions, but legacy injection messages in memory can still impact the agent or other modules, posing long-term risks for LLM agentic systems.
Appendix B Detailed Results on AgentDojo
Table 2: Utility on the AgentDojo benchmark without attack (%)
Model Method Overall Banking Slack Travel Workspace
GPT-4o-mini ReAct 63.55 50.00 66.70 55.00 82.50
DRIFT 58.48 50.00 71.43 50.00 62.50
Claude-3.5-sonnet ReAct 78.25 75.00 90.48 65.00 82.50
DRIFT 75.86 75.00 80.95 65.00 82.50
Claude-3-haiku ReAct 39.97 37.50 52.38 35.00 35.00
DRIFT 37.90 43.75 42.86 25.00 40.00
GPT-4o ReAct 70.86 75.00 80.95 65.00 62.50
DRIFT 73.05 81.25 80.95 65.00 65.00
Qwen2.5 ReAct 26.58 37.50 23.81 10.00 35.00
DRIFT 32.20 50.00 23.81 20.00 35.00
Table 3: Utility on the AgentDojo benchmark under attack (%)
Model Method Overall Banking Slack Travel Workspace
GPT-4o-mini ReAct 48.27 38.19 48.57 47.14 59.17
DRIFT 47.91 40.97 47.62 42.86 60.18
Claude-3.5-sonnet ReAct 52.80 60.42 59.05 47.14 44.58
DRIFT 64.64 71.53 66.67 56.43 63.93
Claude-3-haiku ReAct 33.64 40.28 33.33 29.29 31.67
DRIFT 33.68 32.64 41.90 33.57 26.61
GPT-4o ReAct 55.43 69.44 63.81 64.29 24.17
DRIFT 62.28 72.92 59.05 56.43 60.71
Qwen2.5 ReAct 19.12 24.31 20.00 7.86 24.30
DRIFT 22.20 27.08 19.05 10.71 31.95
Table 4: ASR on the AgentDojo benchmark under attack (%)
Model Method Overall Banking Slack Travel Workspace
GPT-4o-mini ReAct 30.67 34.03 57.14 13.57 17.92
DRIFT 1.29 2.08 0.95 1.43 0.71
Claude-3.5-sonnet ReAct 37.11 8.33 68.57 37.86 33.67
DRIFT 4.41 1.39 0.00 0.00 16.25
Claude-3-haiku ReAct 11.15 11.81 27.62 1.43 3.75
DRIFT 3.03 2.78 3.81 2.86 2.68
GPT-4o ReAct 51.68 62.50 92.38 11.43 40.42
DRIFT 1.53 4.86 0.00 0.00 1.25
Qwen2.5 ReAct 15.11 27.78 24.76 7.86 0.03
DRIFT 0.00 0.00 0.00 0.00 0.00
Appendix C Prompts
Figure 7: The Prompt of constraints Generation.
Figure 8: The Prompt of Privilege Assignment.
Figure 9: The Prompt of Intent Alignment Validation.
Figure 10: The Prompt of Prompt Injection Detection.
Figure 11: The Prompt of Planning Sampling.
Figure 12: The Prompt of Injection Sampling.